OpenStudy (anonymous):

Simpilfy the radical expression -4√160 A.-4√80 B. -4√16 C.-16√10 D.√10

OpenStudy (anonymous):

\[\large \sqrt{160} = \sqrt{4 \times 4 \times 10}\]

OpenStudy (anonymous):

i got b